a body is thrown upwards with velocity of 100m/s and it travel 5m in the last second of its motion.if the same body is thrown upwards with a velocity of 200 m/s find the distance travelled by it in the last second of its journey

Grade:10

1 Answers

Arun
6 years ago
Dear Ishan
 
For case2 there is no effect of case 1.
Now
V² = u² - 2gS
0 = 200² - 2* 10 * S
40000/20 = S
S = 2000 metres
V= u + g*t
0 = 200 - 10 * t
t = 20 sec
In last second distance,
S20 = u + (1/2) g *(2t-1)
 = 200 - (1/2)*10* 39
 = 200 - 195
 = 5 metres
Regards
Arun (askIITians forum expert)